# Crucial-Diff: A Unified Diffusion Model for Crucial Image and Annotation Synthesis in Data-scarce Scenarios

### v1

@@ IV-C Qualitative Comparison
TABLE I: Comparison with other generation methods in MVTec, where Trad. denotes traditional augmentation methods, Gen. denotes generative model based methods, Uni. denotes unified methods and Ano.Diff. denotes AnomalyDiffusion [11]. 1,000 images are generated by each method for training a U-Net as a downstream model. Here, metric Time refers to the total training hours for all targets and metric AP and F1-MAX use pixel-level results.
@@
Trad. Rotate - 1.87 0.16 - 57.55 57.01
Crop-Paste[39] - 1.51 0.14 - 74.08 71.92
Gen. DFMGAN[17] 289.50 1.72 0.20 0.08 63.56 64.49
Ano.Diff.[11] 172.43 1.80 0.32 0.13 82.23 77.07
Uni. Ours 61.15 1.84 0.33 0.07 83.63 78.12
TABLE II: Comparison with other generation methods in polyp dataset. Trad. denotes traditional augmentation methods, Gen. denotes generative model based methods, Uni. denotes unified methods. Metric Time is total training hours, while mDice and mIoU results are the average across test sets. Each method generates 1,450 images to train a SegFormer [59] as downstream model.
[…]
Trad. Rotate - 3.74 0.27 - 81.56 74.03
Crop-Paste[39] - 3.52 0.27 - 80.94 73.16
Gen. SinGAN-Seg[60] 873.75 3.55 0.54 0.07 79.46 71.68
ArSDM[19] 4818.33 3.00 0.56 0.09 82.36 75.18
Uni. Ours 61.15 3.59 0.59 0.05 87.69 81.64
@@ IV-D1 Comparison with Generation Methods
As shown in Table I, our method achieves an IS of 1.84 on the MVTec dataset, significantly surpassing the previous state-of-the-art, AnomalyDiffusion [11], which has an IS of 1.80. The rotational method has the highest IS value because its synthetic output is consistent with the distribution of the original real data, which has a high degree of fidelity. Additionally, Crucial-Diff significantly improves KID value, reducing it to 0.07. The IC-LPIPS score also rises to 0.33, validating that Crucial-Diff can effectively optimise diversity and fidelity simultaneously.
@@
The outstanding performance of Crucial-Diff is further validated on the polyp dataset, as displayed in Table II. Our method achieves an IS of 3.59, a KID of 0.05 and an IC-LPIPS score of 0.59. These results show that Crucial-Diff generalizes well across medical domains, generating crucial images that effectively improve downstream model performance while significantly reducing training time.
@@
TABLE III: Comparison of best anomaly detection performance (pixel-level / image-level) between previous methods and a U-Net model trained on data generated by our method on MVTec dataset. Methods marked with * indicate our reproduced results.
Methods AP $\uparrow$ F1-MAX $\uparrow$
@@
IV-D2 Comparison with Downstream Task Methods
To evaluate the effectiveness of our Crucial-Diff framework in improving downstream task performance, we compare a U-Net model trained on data generated by our method against state-of-the-art methods. In Table III, our Crucial-Diff achieves pixel-level AP of 84.2% and F1-MAX of 78.4%, outperforming both unsupervised and supervised SOTA methods. Notably, compared to PRN [66], which has AP of 78.6% and F1-MAX of 62.3%, our approach significantly improves detecting subtle anomalies. As for the polyp dataset, our Crucial-Diff achieves mDice of 87.9% and mIoU of 81.9%, surpassing the previous method CTNet [67] by 1% in mIoU and 1.2% in mDice, as shown in Table IV. Empirical results demonstrate Crucial-Diff’s consistent performance gains across industrial and medical domains, confirming its task-agnostic generalization capability for diverse downstream applications.
TABLE IV: Segmentation comparisons (mDice / mIoU) between previous methods and a SegFormer [59] model trained on data generated by our method on polyp datasets. Methods marked in bold and underline indicate the highest and second-highest scores.
@@ IV-E1 Effect of Different Components
To assess the contributions of each module in Crucial-Diff, we compare three variants: 1) SAFE Only, which uses only the SAFE module to generate easy samples; 2) WASM Only, which relies solely on the WASM module; and 3) SAFE + WASM (Ours), combining both modules to generate crucial samples. Table V demonstrates the impact of image training on downstream tasks generated under different settings. Only with SAFE module, AP drops by 5.80% and F1-MAX by 4.62%, which means that using only simple data does not allow the downstream model to detect the target better. Only with SRAM module reduces AP by 5.06% and F1-MAX by 1.49%, as it fails to generate target images, forcing the downstream model to learn from non-target samples and misidentify non-target regions. In contrast, our method achieves the highest AP and F1-MAX, demonstrating the necessity of each module.
TABLE V: Ablation study on main components contributions in Crucial-Diff. The MVTec uses pixel-level AP and F1-MAX. The polyp dataset uses the average results of mIoU and mDice.
@@
TABLE VI: Ablation Study of the generated image quality on MVTec for different number of selected text embeddings value $\Gamma$.
@@
TABLE VII: Ablation study on the generated image quality using MVTec dataset for different values of trade-off hyperparameter $\lambda$.
@@
IV-E2 Number of Target Text Embeddings
To investigate the impact of the number of hierarchical features extracted from the CLIP image encoder, we conduct an ablation study on the number of target text embeddings $\Gamma$. The quantitative results are presented in Table VI. As $\Gamma$ increases, the generation performance improves until reaching an optimal configuration at 5, where the synthetic images achieve the superior IS and KID. Beyond $\Gamma=5$, the improvements become marginal, leading us to select 5 as the optimal number of layers. This configuration effectively balances low-level details and high-level semantics, facilitating the synthesis of diverse and realistic images.
IV-E3 Value of Trade-off Hyperparameter $\lambda$
Our ablation study on the trade-off parameter $\lambda$ demonstrates the robustness of the proposed method across different weightings (0.1-0.9) on the MVTec dataset. As shown in Table VII, both IS and KID remain stable, indicating minimal sensitivity to parameter changes. The consistent KID scores particularly highlight our network’s adaptive capability to maintain distribution alignment regardless of $\lambda$ values. This is because $\lambda$ is integrated into the network training process rather than directly controlling the output results. This design ensures stability in the early stages of training while making the model insensitive to specific values of $\lambda$, thereby making the model more robust.

### v2

@@ IV-C2 Distribution Analysis of Synthetic Anomalies
IV-C3 Visualization of Segmentation Results
To further explore the characteristics of crucial samples, we provide a visualization of segmentation results. Fig. 6(a) displays a representative critical sample synthesized by our method, (b) presents the segmentation output of a downstream model trained on only one-third of the original dataset and (c) shows the result from a downstream model trained on 1000 critical samples generated by our approach, evaluated on the same sample. The results demonstrate that training with critical samples enables the model to retain high segmentation sensitivity, particularly toward subtle anomalous features that are often overlooked. By incorporating challenging features designed to target under-learned regions of the decision boundary, critical samples effectively expand the coverage of hard examples beyond what conventional training data provide. Consequently, this approach leads to substantial improvements in the model’s generalization capacity and robustness when deployed in complex and variable environments.
Figure 5: T-SNE visualization of different targets. The pink triangle represents a real image with target, the green circle represents a real image without target, the yellow pentagram represents an easy synthetic image with target, and the purple diamond represents a crucial synthetic image with target.
Figure 6: Visualization comparison of segmentation results on different training data. (a) shows samples generated by our method. (b) shows the segmentation results of a downstream model trained on one-third of the original dataset, while (c) shows the segmentation results of a downstream model augmented with our 1000 key samples.
TABLE I: Comparison with other generation methods in MVTec and Real-IAD, where Trad. denotes traditional augmentation methods, Gen. denotes generative model based methods, Uni. denotes unified methods and Ano.Diff. denotes AnomalyDiffusion [15]. Metric Time refers to the total training hour on a single dataset, while $\dagger$ represents the training time across all datasets.
Methods MVTec Real-IAD
@@ IV-C3 Visualization of Segmentation Results
Trad. Rotate - 1.87 0.16 - - 2.24 0.08 -
Crop-Paste[45] - 1.51 0.14 - - 2.05 0.04 -
Gen. Textual Inversion[35] 113.49 1.66 0.24 0.07 224.62 1.31 0.03 0.11
DFMGAN[23] 289.50 1.72 0.20 0.08 518.82 1.79 0.13 0.24
Ano.Diff.[15] 172.43 1.80 0.32 0.13 206.98 1.90 0.04 0.10
Uni. Ours 61.15 1.84 0.33 0.07 61.15 1.92 0.05 0.10
@@ IV-D1 Comparison with Generation Methods
As summarized in Table I, our proposed method achieves an IS of 1.84 on the MVTec dataset, surpassing AnomalyDiffusion [15] at 1.80 and Textual Inversion [35] at 1.66. Although the rotational baseline achieves the highest IS due to its close distributional alignment with the real training data, our method delivers more comprehensive performance across diverse evaluation metrics. Furthermore, Crucial-Diff reduces the KID to 0.07 and elevates the IC-LPIPS to 0.33, validating that Crucial-Diff can effectively optimise diversity and fidelity simultaneously. While DFMGAN achieves a high IC-LPIPS on Real-IAD by generating distorted and unrealistic images, our method maintains a better balance between diversity and semantic authenticity.
TABLE II: Comparison with other generation methods in five polyp datasets (CLINICDB, ETIS, ENDOSCENE, KVASIR, COLONDB). Trad. denotes traditional augmentation methods, Gen. denotes generative model based methods, Uni. denotes unified methods. Metric Time refers to the total training hour, while $\dagger$ represents the training time across all datasets.
[…]
Trad. Rotate - 3.74 0.27 -
Crop-Paste[45] - 3.52 0.27 -
Gen. Textual Inversion[35] 16.88 3.54 0.13 0.06
SinGAN-Seg[66] 873.75 3.55 0.54 0.07
ArSDM[18] 4818.33 3.00 0.56 0.09
Uni. Ours 61.15 3.59 0.59 0.05
TABLE III: Comparison of mask annotation accuracy across Methods. Quantitative evaluation of synthetic mask quality is performed by computing the IoU between expert-annotated ground truth masks and those generated by different methods. Higher IoU values indicate superior alignment and annotation accuracy.
   Methods    MVTec    Real-IAD    Polyp
   Ano.Diff.    36.73%    30.99%    -
   ArSDM    -    -    83.32%
   Ours    63.48%    44.35%    86.71%
@@
Crucial-Diff also exhibits consistently strong performance on the polyp dataset, as displayed in Table II. Our method achieves an IS of 3.59, a KID of 0.05 and an IC-LPIPS score of 0.59. Notably, the minimal training time of Textual Inversion [35] can be attributed to the fact that it only learns a single target feature, while our method adopts comprehensive training covering all three datasets. These results show that Crucial-Diff generalizes well across medical domains, generating crucial images that effectively improve downstream model performance while significantly reducing training time.
Furthermore, to validate the accuracy of synthetic targets and their segmentation masks, we conduct a systematic evaluation across three public datasets, where ten generated images per category are randomly selected and annotated in detail by domain experts. The IoU is used as the evaluation metric to measure the accuracy of the generated results by calculating the similarity between the generated mask and the expert-drawn mask. As summarized in Table III, our approach consistently outperforms state-of-the-art methods such as AnomalyDiffusion [15] and ArSDM [18] across all datasets, achieving higher IoU scores. These results confirm the high annotation reliability of our generated samples, demonstrating its suitability for training robust and generalizable downstream models.
IV-D2 Comparison of Synthetic Data Effectiveness
In industrial anomaly detection, this study systematically evaluates the proposed synthetic data generation method using both U-Net and PRN [67] architectures on the MVTec dataset. The Real-IAD dataset is excluded from the comparative study since most generative methods exhibit limited generation quality on this benchmark, making downstream performance comparisons less meaningful. As summarized in Table IV, the proposed approach consistently enhances performance across both models. On MVTec, U-Net trained with the synthetic samples achieves 83.6% AP and 78.1% F1-MAX, surpassing all competing methods. The PRN [67] method also shows clear gains when augmented with the synthetic data, demonstrating robust cross-architectural generalization.
For medical image segmentation, all experiments are conducted on a polyp dataset comprising five public datasets. As shown in Table V, SegFormer trained with our synthetic data achieves 87.7% mDice and 81.6% mIoU, while CTNet [68] augmented with our data reaches 88.3% mDice and 82.1% mIoU. Both models outperform those trained on data generated by previous methods, further demonstrating the effectiveness of our synthetic data generation approach.
TABLE IV: Performance of different downstream models trained on synthetic datasets generated by various methods. Each generation method produced 1,000 images on MVTec and Real-IAD datasets. Evaluation using pixel-level AP and F1-MAX metrics. The highest score is highlighted in bold.
[…]
   U-Net
   + rotate    57.6    57.0
   + Crop-Paste[45]    74.1    71.9
   + Textual Inversion[35]    71.1    66.5
   + DFMGAN[23]    63.6    64.5
   + Ano.Diff.[15]    82.2    77.1
   + Ours    83.6    78.1
   PRN[67]
   + rotate    47.3    50.3
   + Crop-Paste[45]    50.4    51.8
   + Textual Inversion[35]    46.2    48.9
   + DFMGAN[23]    39.3    43.2
   + Ano.Diff.[15]    58.6    53.2
   + Ours    63.7    62.0
TABLE V: Performance of different downstream models trained on synthetic datasets generated by various methods. Each generation method produced 1,450 images on five datasets (CLINICDB, ETIS, ENDOSCENE, KVASIR, COLONDB). Evaluation using average mDICE and mIOU metrics across five datasets. The highest score is highlighted in bold.
[…]
   Segformer
   + rotate    81.6    74.0
   + Crop-Paste[45]    80.9    73.2
   + Textual Inversion[35]    67.8    59.2
   + SinGAN-Seg[66]    79.5    71.7
   + ArSDM[18]    82.4    75.2
   + Ours    87.7    81.6
   CTNet[68]
   + rotation    87.0    80.4
   + Crop-Paste[45]    86.7    80.2
   + Textual Inversion[35]    68.5    63.6
   + SinGAN-Seg[66]    68.4    78.5
   + ArSDM[18]    87.2    80.7
   + Ours    88.3    82.1
IV-D3 Comparison with Downstream Task Methods
To evaluate the effectiveness of our Crucial-Diff framework in improving downstream task performance, we compare a U-Net model trained on data generated by our method against state-of-the-art methods. In Table VI, our Crucial-Diff achieves pixel-level AP of 84.2% and F1-MAX of 78.4%, outperforming both unsupervised and supervised SOTA methods. Notably, compared to PRN [67], which has AP of 78.6% and F1-MAX of 62.3%, our approach significantly improves detecting subtle anomalies. As for the polyp dataset, our Crucial-Diff achieves mDice of 87.9% and mIoU of 81.9%, surpassing the previous method CTNet [69] by 1% in mIoU and 1.2% in mDice, as shown in Table VII. Empirical results demonstrate Crucial-Diff’s consistent performance gains across industrial and medical domains, confirming its task-agnostic generalization capability for diverse downstream applications.
@@ IV-D3 Comparison with Downstream Task Methods
TABLE VI: Comparison of best anomaly detection performance (pixel-level / image-level) between previous methods and a U-Net model trained on data generated by our method on MVTec dataset. Methods marked with * indicate our reproduced and marked in bold and underline indicate the highest and second-highest scores.
[…]
@@
TABLE VII: Segmentation comparisons (mDice / mIoU) between previous methods and a SegFormer [75] model trained on data generated by our method on polyp datasets. Methods marked in bold and underline indicate the highest and second-highest scores.
@@ IV-E1 Effect of Different Components
To assess the contributions of each module in Crucial-Diff, we compare three variants: 1) SAFE Only, which uses only the SAFE module for generating samples; 2) WASM Only, which relies exclusively on the WASM module without initialization from the SAFE module; and 3) SAFE + WASM (Ours), combining both modules to generate crucial samples. Table VIII demonstrates the impact of image training on downstream tasks generated under different settings. Only with SAFE module, AP drops by 5.80% and F1-MAX by 4.62%, which means that using only simple data does not allow the downstream model to detect the target better. Only with SRAM module reduces AP by 5.06% and F1-MAX by 1.49%, as it fails to generate target images, forcing the downstream model to learn from non-target samples and misidentify non-target regions. In contrast, our method achieves the highest AP and F1-MAX, demonstrating the necessity of each module.
TABLE VIII: Ablation study on main components contributions in Crucial-Diff. The MVTec uses pixel-level AP and F1-MAX. The polyp dataset uses the average results of mIoU and mDice.
@@
TABLE IX: Ablation study of the generated image quality on MVTec for different number of selected text embeddings value $\Gamma$.
@@
TABLE X: Ablation of fusion strategies on MVTec. Comparative analysis of three paradigms: 1) feature level addition, 2) KV concatenation, and 3) our output integration approach.
[…]
   1)    1.57    0.23    0.07
   2)    1.59    0.24    0.06
   3)    1.84    0.33    0.07
IV-E2 Number of Target Text Embeddings
To investigate the impact of the number of hierarchical features extracted from the CLIP image encoder, we conduct an ablation study on the number of target text embeddings $\Gamma$. The quantitative results are presented in Table IX. As $\Gamma$ increases, the generation performance improves until reaching an optimal configuration at 5, where the synthetic images achieve the superior IS and KID. Beyond $\Gamma=5$, the improvements become marginal, leading us to select 5 as the optimal number of layers. This configuration effectively balances low-level details and high-level semantics, facilitating the synthesis of diverse and realistic images.
IV-E3 Fusion Strategies of cross attention
To validate the design of the fusion operation within the cross-attention mechanism, we conduct a comparative ablation: 1) feature level addition: $\mathbf{V}_{t}\leftarrow\lambda\cdot\mathbf{V}^{c}_{t}+(1-\lambda)\cdot\mathbf{V}^{u}_{t}$ and $\mathbf{K}_{t}\leftarrow\lambda\cdot\mathbf{K}^{c}_{t}+(1-\lambda)\cdot\mathbf{K}^{u}_{t}$; 2) KV concatenation: $Attn\leftarrow Attn(\mathbf{Q}_{t},[\mathbf{K}_{t}^{c};\mathbf{K}_{t}^{u}],[\mathbf{V}_{t}^{c};\mathbf{V}_{t}^{u}])$; and 3) our output integration approach. Table X demonstrate the superior performance of our approach. Our method improves IS by 0.27 over feature addition and 0.25 over KV concatenation. It also increases IC-LPIPS by 0.10 and 0.09 over the two baselines, respectively. The strategy we propose enables stable and controllable integration of deceptive features and semantic constraints, whereas other fusion methods lead to training instability issues.
IV-E4 Value of Trade-off Hyperparameter $\lambda$
Our ablation study on the trade-off parameter $\lambda$ demonstrates the robustness of the proposed method across different weightings (0.1-0.9) on the MVTec dataset. As shown in Table XI, both IS and KID remain stable, indicating minimal sensitivity to parameter changes. The consistent KID scores particularly highlight our network’s adaptive capability to maintain distribution alignment regardless of $\lambda$ values. This is because $\lambda$ is integrated into the network training process rather than directly controlling the output results. This design ensures stability in the early stages of training while making the model insensitive to specific values of $\lambda$, thereby making the model more robust.
TABLE XI: Ablation study on the generated image quality using MVTec dataset for different values of trade-off hyperparameter $\lambda$.